# ClaimVer: Explainable Claim-Level Verification and Evidence Attribution of Text Through Knowledge Graphs

### v1

@@ 6 Discussion
While there are several advantages associated with using KGs, we also acknowledge the presence of known issues, such as knowledge coverage and the efforts required to keep these sources up-to-date. For our solution, we assume that the KG is up-to-date and possesses adequate coverage. However, this may not always be the case, and thus the most suitable technique should be adopted after considering the specific requirements of a particular use case. Another point to consider is that the proposed method does not provide traditional citations to articles, although it may be possible to retrieve that information from the KG, if information source mapping has been properly maintained.
@@ 7 Conclusion
In this paper, we present ClaimVer, a framework that facilitates text verification and evidence attribution at the claim level by leveraging information present in KGs. We have prioritized human-centric design principles to make the framework more informative, intuitive, and user-friendly. Additionally, our methodology incorporates design choices that ensure open access, sustainability, and reliability.
ClaimVer presents several advantages, as outlined below:
  1. 1.
    Human-centric design: In addition to its primary functions of text verification and evidence attribution, the system generates considerable information conducive to user awareness. This information serves to educate users and enhance their trust in the automated system.
  2. 2.
    Finer Granularity: Perform validation at the claim level, enabling localization of hallucinations, or false claims.
  3. 3.
    Enhanced Coverage: Eliminate one-to-one mapping between input and reference text, allowing for layered interpretation, and handling of distributed information.
  4. 4.
    Domain Adaptability: Flexibility in adapting to new domains by switching to a more suitable KG.
  5. 5.
    Maintenance Efficiency: Simplified auditing and updation of the knowledge base, ensuring the data remains current and accurate.

### v2

@@ 7 Conclusion
In this paper, we present ClaimVer, a framework for text verification and evidence attribution at the claim level by leveraging information present in KGs. In contrast to other methods, ClaimVer eliminates the one-to-one mapping between input and reference text, allowing for layered interpretation and handling of distributed information. In addition to these primary functions, ClaimVer incorporates human-centric design principles by offering clear, concise explanations for each claim prediction—an important characteristic for building user trust and enhancing usability. Furthermore, we introduce an attribution score, which enhances its applicability across a wide range of downstream tasks. Finally, we share ClaimVer fine-tuned LLMs to facilitate further exploration of this research direction.
8 Limitations
Limitations of LLMs for Fact Verification. Like most ML models, LLMs are prone to erroneous predictions, which is particularly concerning in sensitive applications such as handling misinformation. Despite this, they remain the most performant techniques for fact verification and related tasks like NLI Yue et al. (2023); Wang et al. (2021b). Therefore, while it is reasonable to use the best option available, fact verification systems relying on LLMs should be utilized with caution and necessary validations.
Limitations of Knowledge Graph. While there are several advantages associated with using KGs, we also acknowledge the presence of known issues, such as knowledge coverage and the efforts required to keep these sources up-to-date. For our solution, we assume that the KG is up-to-date and possesses adequate coverage. However, this may not always be the case, and thus the most suitable technique should be adopted after considering the specific requirements of a particular use case. Another point to consider is that the proposed method does not provide traditional citations to articles, although it may be possible to retrieve that information from the KG, if information source mapping has been properly maintained.
Variations in Claim Decomposition. Decomposing text into multiple claims is a complex linguistic task that often results in multiple valid decompositions. Although this may not impact usability if the prediction, rationale, and text spans are comprehensible and supported by facts from the KG, it poses a challenge for evaluating model performance. One potential approach is to operate at the token level instead of the span level, but this would significantly complicate the problem space. Additionally, token-level verification and attribution would require substantially higher compute resources, necessitating further studies to assess their value and impact on system usability and reliability.
LLM Reasoning Errors. Previous works have demonstrated that using LLM reasoning for tasks like fact verification, evidence attribution, and NLI can yield impressive results, surpassing alternative approaches Yue et al. (2023); Wang et al. (2021b). However, LLM reasoning can still be flawed. In our work, we impose validations to minimize LLM mistakes by performing membership checks for supporting triplets and string matching for text spans. However, validating reasoning remains an open problem with ongoing research efforts.
Fine-tuning Dataset Limitations. To build the fine-tuning dataset, we utilized GPT-4 with two detailed sequential prompts designed in accordance with OpenAI’s recommendations OpenAI (2024) and previous works Nori et al. (2023). Despite employing techniques like few-shot prompting with state-of-the-art LLMs, we still observe mistakes, indicating the complexity of this problem. To address this, we conducted manual checks to minimize errors and share the dataset with the research community for further improvement.
9 Ethical Concerns
Our work presents a scalable and interpretable framework for fact-checking textual claims. To promote the exploration of this important problem space, we fine-tune and share open-source LLMs that are well-aligned to the framework’s objective function. While the models we provide perform better than the publicly available base models for our specific task, they still share similar weaknesses. To address this as best as we can, we have incorporated and described ways to mitigate these issues to the extent possible. We believe the benefits of this work outweigh any potential risks.
Appendix A Appendix
A.1 Training Details
In this section, we present the training parameters used for fine-tuning each model, along with their corresponding loss plots. All models converged after two epochs, achieving ROUGE-L Lin (2004) scores greater than 0.658, with the highest model reaching 0.719.
Figure 4: Fine-tuning loss plots for Llama3-8B-Chat.
Parameter Value
Base Model meta-llama/Meta-Llama-3-8B-Instruct
ROUGE-L TBD
ROUGE-1 TBD
Fine-Tuning Type LoRA
LoRA Alpha 16
LoRA Rank 8
Cutoff Length 4096
Gradient Accumulation Steps 8
Learning Rate 5.0e-05
LR Scheduler Type Cosine
Number of Training Epochs 2.0
Optimizer AdamW
Quantization Bit 4
Table 5: Fine-tuning Parameters for Llama3-8B-Chat
Figure 5: Fine-tuning loss plots for Mistral-7B-v0.3-Chat.
Parameter Value
Base Model mistralai/Mistral-7B-Instruct-v0.3
ROUGE-L 0.694
ROUGE-1 0.719
Fine-Tuning Type LoRA
LoRA Alpha 16
LoRA Rank 8
Cutoff Length 4096
Gradient Accumulation Steps 8
Learning Rate 5.0e-05
LR Scheduler Type Cosine
Number of Training Epochs 2.0
Optimizer AdamW
Quantization Bit 4
Table 6: Fine-tuning Parameters for Mistral-7B-v0.3-Chat
Figure 6: Fine-tuning loss plots for Phi-3-mini-4k-Chat.
Parameter Value
Base Model microsoft/Phi-3-mini-4k-instruct
ROUGE-L 0.658
ROUGE-1 0.685
Fine-Tuning Type LoRA
LoRA Alpha 16
LoRA Rank 8
Cutoff Length 4096
Gradient Accumulation Steps 8
Learning Rate 5.0e-05
LR Scheduler Type Cosine
Number of Training Epochs 2.0
Optimizer AdamW
Quantization Bit 4
Table 7: Fine-tuning Parameters for Phi-3-mini-4k-Chat
Figure 7: Fine-tuning loss plots for SOLAR-10.7B-Chat.
Parameter Value
Base Model upstage/SOLAR-10.7B-Instruct-v1.0
ROUGE-L 0.689
ROUGE-1 0.714
Fine-Tuning Type LoRA
LoRA Alpha 16
LoRA Rank 8
Cutoff Length 4096
Gradient Accumulation Steps 8
Learning Rate 5.0e-05
LR Scheduler Type Cosine
Number of Training Epochs 2.0
Optimizer AdamW
Quantization Bit 4
Table 8: Fine-tuning Parameters for SOLAR-10.7B-Chat
Figure 8: Fine-tuning loss plots for Vicuna-7B-v1.5-Chat.
Parameter Value
Base Model lmsys/vicuna-7b-v1.5
ROUGE-L 0.676
ROUGE-1 0.700
Fine-Tuning Type LoRA
LoRA Alpha 16
LoRA Rank 8
Cutoff Length 4096
Gradient Accumulation Steps 8
Learning Rate 5.0e-05
LR Scheduler Type Cosine
Number of Training Epochs 2.0
Optimizer AdamW
Quantization Bit 4
Table 9: Fine-tuning Parameters for Vicuna-7B-v1.5-Chat
Figure 9: Fine-tuning loss plots for Zephyr-7B-Beta-Chat.
Parameter Value
Base Model HuggingFaceH4/zephyr-7b-beta
ROUGE-L 0.686
ROUGE-1 0.712
Fine-Tuning Type LoRA
LoRA Alpha 16
LoRA Rank 8
Cutoff Length 4096
Gradient Accumulation Steps 8
Learning Rate 5.0e-05
LR Scheduler Type Cosine
Number of Training Epochs 2.0
Optimizer AdamW
Quantization Bit 4
Table 10: Fine-tuning Parameters for Zephyr-7B-Beta-Chat
Figure 10: Fine-tuning loss plots for Gemma-7B-IT-Chat.
Parameter Value
Base Model google/gemma-7b-it
ROUGE-L 0.678
ROUGE-1 0.703
Fine-Tuning Type LoRA
LoRA Alpha 16
LoRA Rank 8
Cutoff Length 4096
Gradient Accumulation Steps 8
Learning Rate 5.0e-05
LR Scheduler Type Cosine
Number of Training Epochs 2.0
Optimizer AdamW
Quantization Bit 4
Table 11: Fine-tuning Parameters for Gemma-7B-IT-Chat
Figure 11: Fine-tuning loss plots for Gemma-2B-IT-Chat.
Parameter Value
Base Model google/gemma-2b-it
ROUGE-L 0.667
ROUGE-1 0.692
Fine-Tuning Type LoRA
LoRA Alpha 16
LoRA Rank 8
Cutoff Length 4096
Gradient Accumulation Steps 8
Learning Rate 5.0e-05
LR Scheduler Type Cosine
Number of Training Epochs 2.0
Optimizer AdamW
Quantization Bit 4
Table 12: Fine-tuning Parameters for Gemma-2B-IT-Chat
**Text Span Attribution Verification**
**Objective:** Predict whether the text span is "Attributable", "Contradictory", or "Extrapolatory" based on the information provided in the triplets.
**Instructions:**
1. **Read the Full Text:**
- Understand the context and content of the full text string.
2. **Examine the Text Span:**
- Determine the claims made within the text span.
3. **Analyze the Triplets:**
- Evaluate if the triplets support, refute, or neither support nor refute the claims in the text span.
4. **Make Your Prediction:**
- Classify the text span as "Attributable", "Contradictory", or "Extrapolatory" based on your analysis of the triplets.
5. **Provide Rationale:**
- Clearly explain your reasoning for the classification.
**Classification Criteria:**
- **"Attributable"**: The text span is sufficiently supported by the triplet(s). All claims in the text span are directly present in the triplet information.
- **"Contradictory"**: The text span is conclusively refuted by the triplet(s). All claims in the text span are directly contradicted by the triplet information.
- **"Extrapolatory"**: The triplet(s) can neither support nor refute the text span. The information provided is either irrelevant, indirect, or related but not sufficient to support or refute the text span.
**Example:**
**Full Text:** "Albert Einstein is widely recognized as the father of modern physics. He was awarded the Nobel Prize in Physics for his services to Theoretical Physics."
**Text Span:** "He was awarded the Nobel Prize in Physics."
**Triplets:** [("Albert Einstein", "award received", "Nobel Prize in Physics")]
**Sample Evaluation:**
- **Prediction:** "Attributable"
- **Rationale:** "The triplet directly supports the claim that Albert Einstein received the Nobel Prize in Physics."
**Example:**
**Full Text:** "Isaac Newton discovered the element radium."
**Text Span:** "Isaac Newton discovered radium."
**Triplets:** [("Marie Curie", "discovered", "radium")]
**Sample Evaluation:**
- **Prediction:** "Contradictory"
- **Rationale:** "The triplet states that Marie Curie discovered radium, contradicting the claim that Isaac Newton discovered it."
**Example:**
**Full Text:** "The Eiffel Tower is a wrought-iron lattice tower that was opened in 1889."
**Text Span:** "The Eiffel Tower is a wrought-iron lattice tower that was opened in 1889."
**Triplets:** [("Eiffel Tower", "located in", "Paris")]
**Sample Evaluation:**
- **Prediction:** "Extrapolatory"
- **Rationale:** "The triplet states that the Eiffel Tower is located in Paris, which is related but not sufficient to confirm or refute that it was opened in 1889."
**Verification Checklist:**
- [ ] The prediction accurately reflects the relationship between the text span and the triplets.
- [ ] The rationale clearly explains the classification based on the triplets.
- [ ] The explanation is free from irrelevant information.
**Response Format:**
Provide your evaluation in the following JSON format:
- "prediction": "Attributable", "Contradictory", or "Extrapolatory"
- "rationale": "Your comments here"
**Inputs to Evaluate**
**Full text:** "{full_text}"
**Text span:** "{text_span}"
**Triplets:** {triplets}
Figure 12: Prompt for the text span attribution verification task, guiding the model to classify text spans as "Attributable," "Contradictory," or "Extrapolatory" based on the provided triplets. The prompt design incorporates concepts such as few-shot learning, chain-of-thought reasoning Kojima et al. (2022), and tailored prompt engineering OpenAI (2024); Nori et al. (2023)